# Revisit, Extend, and Enhance Hessian-Free Influence Functions

### v1

@@ Abstract
Influence functions serve as crucial tools for assessing sample influence in model interpretation, subset training set selection, noisy label detection, and more. By employing the first-order Taylor extension, influence functions can estimate sample influence without the need for expensive model retraining. However, applying influence functions directly to deep models presents challenges, primarily due to the non-convex nature of the loss function and the large size of model parameters. This difficulty not only makes computing the inverse of the Hessian matrix costly but also renders it non-existent in some cases. Various approaches, including matrix decomposition, have been explored to expedite and approximate the inversion of the Hessian matrix, with the aim of making influence functions applicable to deep models. In this paper, we revisit a specific, albeit naive, yet effective approximation method known as TracIn. This method substitutes the inverse of the Hessian matrix with an identity matrix. We provide deeper insights into why this simple approximation method performs well. Furthermore, we extend its applications beyond measuring model utility to include considerations of fairness and robustness. Finally, we enhance TracIn through an ensemble strategy. To validate its effectiveness, we conduct experiments on synthetic data and extensive evaluations on noisy label detection, sample selection for large language model fine-tuning, and defense against adversarial attacks.
@@ 1 Introduction
Data-centric learning is an emerging topic focusing on enhancing algorithmic performance from the perspective of training data [1] and aims to address data quality issues coupled with learning algorithms. In contrast to model-centric learning, which designs novel algorithms or optimization techniques for performance improvement with fixed training data, data-centric learning operates with a fixed learning algorithm while modifying the training data through trimming, augmenting, or other methods aligned with improving utility. Data-centric learning holds significant potential in many areas such as model interpretation, subset training set selection, data generation, noisy label detection, active learning, domain adaptation, fairness machine learning, and others [2, 3].
@@
As the foundation of data-centric learning, influence estimation and data valuation approaches can be generally categorized into two categories [4]. (a) Retraining-based methods include the classical leave-one-out influence approach [5] retraining models and observing performance changes after removing one (or a group of) training sample(s). While useful as an ideal baseline, even on computation non-extensive models this is computationally untenable. Other representative methods such as model-agnostic Shapley value approaches [6, 7, 8] also suffer similar problems. Computationally efficient approaches such as KNN-Shap [9] tend to make assumptions about the learning model and hence are not directly applicable to other models being considered. (b) Gradient-based methods can be used to approximately estimate influence without expensive overheads of retraining. The seminal work in this category is that of [10], which utilizes a Taylor-series approximation and LiSSA optimization [11] to compute sample influences. However, the limiting assumption is that the model and loss function are convex. Despite debates on the necessity of convexity [12, 13, 14, 15], challenges persist when directly applying gradient-based methods to large models. The size of model parameters complicates calculations, particularly in obtaining the inverse of the Hessian matrix. Efforts, including matrix decomposition techniques [10, 13, 3], aim to expedite and approximate Hessian matrix inversion, therefore enhance the feasibility of influence functions for deep learning models.
@@
Contributions. In this paper, we revisit a particular naive yet aggressive approximation method TracIn [16] by substituting the inverse of the Hessian matrix with an identity matrix, representing it as the Hessian-free influence function, the inner product of the gradient of the validation set and the gradient of the sample under evaluation, which we term Inner Product (IP). We summarize our major contributions as follows:
@@
  • We redefine the TracIn method into IP formulation and delve into the rationale behind this simple yet effective approximation, offering insights into why it performs well in practical scenarios.
@@
  • Expanding on the TracIn framework, we extend its applicability beyond measuring sample influence on model utility to encompass considerations of fairness and robustness.
  • To enhance the generalization of TracIn, we propose IP Ensemble, a novel approach leveraging dropout mechanisms to simulate diverse models. IP Ensemble amalgamates IP scores from these varied models, thus increasing the method’s generalization capabilities.
@@
  • We validate the effectiveness of IP through synthetic data experiments and conduct extensive real-world evaluations using IP Ensemble. These experiments span various applications, including noisy label correction for vision data, data curation aimed at fine-tuning fairer NLP models, and defense strategies against adaptive evasion adversaries.
@@ 2 Related Work
Efficient Influence Estimation. Influence functions serve as crucial tools for estimating the individual valuation of data without requiring model retraining. However, the computation of the inverse of the Hessian matrix poses challenges for large-scale data and models. To address this issue, various approaches have been proposed to simplify or estimate the inverse of the Hessian matrix effectively. A seminal work in this domain is that of [10], which employs a Taylor-series approximation and LiSSA optimization [11] to compute sample influences. Similarly, EKFAC [13] enhances K-FAC by offering a more precise Hessian approximation through efficient eigendecomposition and managing damped inverse Hessian-vector products. More recently, DataInf [3] efficiently computes influence even for large models by replacing the inverse Hessian computation with a readily computable closed-form expression, although their framework may suffer from significant theoretical errors. TracInc [16], a straightforward yet aggressive approximation, substitutes the inverse of the Hessian matrix with an identity matrix, essentially considering gradients directly as a measure of influence. Beyond the conventional influence function that gauges sample influence on the validation set, self-influence [17, 18] computes influence using the training set alone.
@@
Miscellaneous. Several studies have examined the fragility of influence functions in explaining deep learning model predictions. [12] breaks down the difference between influence function estimates and leave-one-out retraining in neural networks into five factors, discovering that while influence estimates may not perfectly align with leave-one-out retraining, they approximate the proximal Bregman response function, offering valuable insights for identifying influential or mislabeled examples. [14] demonstrates that the effectiveness of influence functions in neural networks varies with network architecture, depth, width, parameterization, and regularization, underscoring their fragility in deep learning due to non-convex loss functions. [15] suggests that the instability of current validation procedures, rather than non-convexity or lack of regularization, may be responsible for their unreliability. Other topics include [40] expands influence functions from estimating the effects of removing one point to large groups of training samples; generalized influence functions [41] enhance influence estimation in large-scale models by concentrating on target parameters and addressing computational instability with a robust inverse-Hessian-vector product approximation; multi-stage influence function [42] extends traditional influence functions to monitor the impact of pre-training data on fine-tuned model predictions, facilitating the identification of crucial pre-training examples.
@@ 3 Methods
Revisit. Given a training set $T$$=$$\{z_{i}$$=$$(x_{i},y_{i})\}_{i=1}^{n}$ and a classifier trained using empirical risk minimization by a convex loss function $\ell$, the optimal parameters of the classifier can be obtained by $\hat{\theta}=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(z_{i};\theta)$. To measure the impact of an individual data sample, we can train the model with and without the specific sample and see the performance change. However, the retrain-based approach is expensive for large-scale data and models. To avoid model retraining, influence functions estimate the effect of changing an infinitesimal weight of samples on a validation set $V$$=$$\{z_{j}$$=$$(x_{j},y_{j})\}_{j=1}^{n^{\prime}}$, based on an impact function $f$ evaluating the quantity of interest. If we measure the sample impact on model utility, i.e., the loss on the validation set, by removing this sample from the original training set, the sample influence can be estimated as follows [10]:
@@
The above equation is a Hessian-free approximation, which in essence is the inner product of $\nabla v^{\textrm{util}}$ and the sample gradient $\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$; therefore, we call this method Inner Product (IP) in our paper. In the original paper [16], the authors employ gradient tracing to measure the sample influence, and explain its rationality from calculus; in the following, we link the IP to influence functions and provide our insights of IP on why such a naive approximation works well in practice.
@@
Let $\nabla v^{\textrm{util}}$$=$$\sum\nolimits_{z_{j}\in V}\nabla_{\hat{\theta}}\ell(z_{j};\hat{\theta})^{\top}$, then the influence functions in Eq. (1) and (2) can be reformulated as $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$ and $\nabla v^{\textrm{util}}\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$. It is worth noting that given the well-trained model parameters $\hat{\theta}$, $\nabla v^{\textrm{util}}$ and $\mathbf{H}^{-1}_{\hat{\theta}}$ remain constant. We can visualize the direction of $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ and $\nabla v^{\textrm{util}}$ in Figure 1, with the influence scores roughly representing the angle between the sample gradient and $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ or $\nabla v^{\textrm{util}}$, where $\mathbf{H}^{-1}_{\hat{\theta}}$ induces a rotation of $\alpha$ between these two directions. Remarkably, although the influence scores calculated by Eq. (1) and (2) may differ, they exhibit order-consistency in several scenarios. For instance, if two samples $z_{i}$ and $z_{i^{\prime}}$ have gradients in Region II or IV, then $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ implies $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$. However, this order-consistency does not hold in Regions I and III. Fortunately, the samples in these regions are jointly recognized by $\mathcal{I}^{\textrm{util}}$ and $\mathcal{I}^{\textrm{util}}_{\textrm{IP}}$ as beneficial or detrimental, which has little practical difference. A complete analysis can be found in Appendix A.
@@
In general, IP emerges as a straightforward, efficient, and remarkably effective alternative to influence functions, particularly in convex settings. However, in non-convex scenarios, approximations of the inverse Hessian matrix may introduce significant errors, rendering $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ ineffective, even incorrect. We substantiate this observation empirically in Section 4. Conversely, while $\nabla v^{\textrm{util}}$ in IP may not precisely align with the optimal direction for enhancing model performance, it remains a reliable indicator for distinguishing between detrimental and beneficial samples, even for non-convex models. Intuitively, if a sample gradient aligns with the gradient from the validation set, it suggests that incorporating this sample contributes to enhancing the model’s utility. Similarly, the negative impact of a single sample on prediction can also be mitigated by considering its gradient.
Figure 2: Illustrating our inner product approach on two synthetic datasets and convex/non-convex models. A-C illustrate a 2D linearly separable synthetic dataset with a subset of detrimental samples bearing incorrect labels, trained using a Logistic Regression model for binary classification, and D-F demonstrate the same analysis on a non-linear synthetic half-moon dataset with mislabeled samples, trained using a Multilayer Perceptron neural network. Specifically, A and D depict training sets with class labels 0 (red) and 1 (blue), where detrimental samples are marked with $\times$ and regular samples with $\circ$. B and E show test sets utilized to assess model performance under convex and non-convex conditions. C and F present the calculated influence scores and inner product scores derived by Eq. (1) and (2), respectively. In the linear dataset with the convex model, there is a clear correlation between influence scores and inner product scores, the detrimental samples have both negative influence scores and inner product scores. However, in the non-linear dataset with the non-convex model, the influence scores of detrimental samples appear intermixed; fortunately, the detrimental samples are effectively isolated from inliers through inner product scores.
@@
Enhancement: The simplicity inherent in the IP framework offers opportunities to enhance the generalization of influence functions. In convex cases, the model parameter $\hat{\theta}$ is both optimal and unique. However, in non-convex scenarios, the presence of local minima introduces instability and non-uniqueness into the solution. Typically, ensemble strategies are employed to bolster model generalization. Yet, while employing different models can enhance performance, it also escalates the costs associated with model training and complicates the calculation of influence functions. This arises from the variability in model parameters, necessitating multiple computations of the inverse of each individual Hessian matrix. The introduction of Hessian-free IP circumvents this issue, eliminating the need for costly calculations of Hessian matrices and their inverses. Drawing inspiration from dropout mechanisms, diverse models can be swiftly generated without necessitating model retraining. By computing sample gradients from various models, we propose IP Ensemble that amalgamates IP scores from distinct models. Experiments detailed in Section 5 illustrate the superior performance of the inner product ensemble over other influence function-based methods.
@@ 4 Correctness Verification on Synthetic Data
Here we verify the correctness of our Inner Product (IP) as an influence score surrogate on two synthetic datasets with convex/non-convex models. First, we use Logistic Regression on a linear dataset to demonstrate the effectiveness of IP in a convex optimization scenario. Second, we use a Multi Layer Perceptron on a half-moon dataset to validate the performance of IP in a non-convex optimization scenario. [^1] The detailed characteristics of datasets and models used in this paper can be found in Appendix B. In Figure 2, A and B illustrate the training and test sets of a linearly separable dataset, consisting of 150 and 100 samples, respectively. Notably, the training set contains 10 manually generated noisy samples with incorrect labels. Similarly, D-F illustrate a non-linear separable half-moons dataset, consisting of 200 training samples including 20 noisy samples with incorrect labels, and 100 test samples. C and F display the influence score and inner product score for each training sample, calculated according to Eq. (1) and (2), respectively. Beyond model utility, we also verify the correctness of IP on model fairness and robustness in Appendix C.1.
Figure 3: Directions of gradients of the validation set and detrimental samples of the datasets in Figure 2. In the linear case, $\alpha$ is $0.66\degree$, and we draw a larger angle for better visualization. For the non-linear case, we use the dimension reduction [45] for visualization, and $\alpha$ is $94.36\degree$.
@@
However, the limitations of influence scores become apparent in the context of non-convex models, as illustrated in Figure 2F. Here, the influence scores of detrimental samples are intermingled with those of normal ones, where $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ is not the optimal direction due to the inaccuracies in approximating the Hessian matrix. Fortunately, the IP score effectively isolates detrimental samples from inliers. Even for this non-linear dataset, $\nabla v^{\textrm{util}}$ remains a useful indicator for discerning detrimental samples from beneficial ones, as all detrimental samples exhibit obtuse angles with $\nabla v^{\textrm{util}}$.

### v2

@@ Abstract
Influence functions serve as crucial tools for assessing sample influence. By employing the first-order Taylor extension, sample influence can be estimated without the need for expensive model retraining. However, applying influence functions directly to deep models presents challenges, primarily due to the non-convex nature of the loss function and the large size of model parameters. This difficulty not only makes computing the inverse of the Hessian matrix costly but also renders it non-existent in some cases. Various approaches, including matrix decomposition, have been explored to expedite and approximate the inversion of the Hessian matrix, with the aim of making influence functions applicable to deep models. In this paper, we revisit a specific, albeit naive, yet effective approximation method known as TracIn, and simplify it further, introducing the name Inner Product (IP). This method substitutes the inverse of the Hessian matrix with an identity matrix. We offer deeper insights into why this straightforward approximation method is effective. Furthermore, we extend its applications beyond measuring model utility to include considerations of fairness and robustness. Finally, we enhance IP through an ensemble strategy. To validate its effectiveness, we conduct experiments on synthetic data and extensive evaluations on noisy label detection, sample selection for large language model fine-tuning, and defense against adversarial attacks.
@@ 1 Introduction
Data-centric learning is a growing research field that focuses on enhancing machine learning model performance by refining the quality and characteristics of training data [42]. In contrast to model-centric approaches, which prioritize improving algorithms or optimization techniques without altering the dataset, data-centric learning involves adjusting the dataset itself—through methods like trimming, relabeling, and reweighting—while keeping the learning algorithm fixed. This approach plays a vital role in areas such as model interpretability, selecting training subsets, generating synthetic data, detecting noisy labels, improving active learning, and promoting fairness learning [11, 31].
@@
Sample influence estimation, as the foundation of data-centric learning, can be generally categorized into two categories [20]. (a) Retraining-based methods assess the sample influence by retraining the model with and without a specific sample and checking the performance change, which include the classical leave-one-out influence approach [13] and Shapley value approaches [18, 24, 30, 23]. (b) Gradient-based methods estimate influence without expensive overheads of retraining, known as influence functions. The seminal work in this category is that of [26], which utilizes a Taylor-series approximation and LiSSA optimization [1] to compute sample influences. However, the limiting assumption is that the model and loss function are convex. Despite debates on the necessity of convexity [2, 19, 4, 17], challenges persist when directly applying gradient-based methods to large models. The size of model parameters complicates calculations, particularly in obtaining the inverse of the Hessian matrix. Efforts, including matrix decomposition techniques [26, 19, 31], aim to expedite and approximate Hessian matrix inversion, therefore enhance the feasibility of influence functions for deep learning models.
@@
Contributions. In this paper, we focus on the influence function category and revisit a specific naive yet aggressive approximation method, TracIn [46]. This method substitutes the inverse of the Hessian matrix with an identity matrix, representing it as a Hessian-free influence function—the inner product of the gradient of the validation set and a sample to be assessed, which we refer to as Inner Product (IP). We summarize our major contributions as follows:
@@
  • We simplify the checkpoints of TracIn, define it into IP formulation, and delve into the rationale behind this simple approximation, offering insights into why it performs well in practice.
@@
  • Expanding our IP framework, we extend its applicability beyond measuring sample influence on model utility to encompass considerations of fairness and robustness.
  • To improve generalization, we introduce IP Ensemble, an innovative approach that employs dropout mechanisms to simulate a diverse array of models. IP Ensemble consolidates IP scores from these different models, thereby strengthening the method’s overall generalization capabilities.
@@
  • By assessing the sample influence, model performance can be enhanced by trimming detrimental samples from the training or fine-tuning set, downweighting them, or even relabeling them. Specifically, we validate the effectiveness of IP through synthetic data experiments and conduct extensive real-world evaluations using IP Ensemble. These experiments span various applications, including noisy label correction for vision data, data curation aimed at fine-tuning fairer NLP models, and defense strategies against adaptive evasion adversaries.
@@ 2 Related Work
Efficient Influence Estimation. Influence functions serve as crucial tools for estimating the individual valuation of data without requiring model retraining. However, the computation of the inverse of the Hessian matrix poses challenges for large-scale data and models. To address this issue, various approaches have been proposed to simplify or estimate the inverse of the Hessian matrix effectively. A seminal work is that of [26], which employs a Taylor-series approximation and LiSSA optimization [1] to compute sample influences. Arnoldi [48] employs the random projection and simplified Hessian matrix for acceleration. EKFAC [19] enhances Kronecker-Factored eigendecomposition for a precise Hessian approximation. More recently, DataInf [31] efficiently computes influence even for large models by replacing the inverse Hessian computation with a readily computable closed-form expression, although their framework may suffer from significant theoretical errors. TracInc [46], a straightforward yet aggressive approximation, substitutes the inverse of the Hessian matrix with an identity matrix, essentially considering gradients directly as a measure of influence. Beyond the conventional influence function that gauges sample influence on the validation set, self-influence [5, 52] computes influence using the training set alone. Moving beyond using a single model checkpoint, GEX [25] leverages a geometric ensemble of multiple checkpoints to approximate influence functions, alleviating the bilinear constraint and non-linear losses. Moreover, TDA [3] also introduces a checkpoint-based segmentation approach, which further enhances the training process by combining the techniques of implicit differentiation and unrolling. This is achieved through the utilization of EKFAC, providing a more efficient and scalable solution for complex models [19].
@@
Miscellaneous. Several studies have examined the fragility of influence functions in explaining deep learning model predictions. [2] discovers that while influence estimates may not perfectly align with leave-one-out retraining, they approximate the proximal Bregman response function, offering valuable insights for identifying influential or mislabeled examples. [4] demonstrates that the effectiveness of influence functions in neural networks varies with network architecture, depth, width, parameterization, and regularization, underscoring their fragility in deep learning due to non-convex loss functions. [17] suggests that the instability of current validation procedures, rather than non-convexity or lack of regularization, may be responsible for their unreliability. [27] expands influence functions from estimating the effects of removing one point to large groups of training samples; [38] enhance influence estimation in large-scale models by concentrating on target parameters and addressing computational instability with a robust inverse-Hessian-vector product approximation;  [9] extend traditional influence functions to monitor the impact of pre-training data on fine-tuned model predictions, facilitating the identification of crucial pre-training examples.
@@ 3 Methods
Revisit and Simplify. Given a training set $T$$=$$\{z_{i}$$=$$(x_{i},y_{i})\}_{i=1}^{n}$ and a classifier with empirical risk minimization by a convex loss function $\ell$, the optimal parameters of the classifier can be obtained by $\hat{\theta}=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(z_{i};\theta)$. To measure the influence of an individual data sample, we can train the model with and without the specific sample and see the performance change. However, the retrain-based approach is expensive for large-scale data and models. To avoid model retraining, influence functions estimate the effect of changing an infinitesimal weight of samples on a validation set $V$$=$$\{z_{j}$$=$$(x_{j},y_{j})\}_{j=1}^{n^{\prime}}$, based on an impact function $f$ evaluating the quantity of interest. Considering the sample impact on model utility, i.e., the loss on the validation set, by removing this sample from the original training set, the sample influence can be estimated as follows [26]:
@@
Figure 1: Illustration of $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ and $\nabla v^{\textrm{util}}$.
The above equation is the inner product of $\nabla v^{\textrm{util}}$ and the sample gradient $\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$; therefore, we call this method Inner Product (IP). Note that TracIn incorporates multiple checkpoints to record model parameters throughout the optimization process, whereas IP only takes into account the final or converged model. We believe that sample influence should be assessed using a fixed model; simply summing sample influences at different stages fails to capture the dynamics of model optimization. The converged model can significantly diverge from its earlier stages, making the sample influences derived from initial checkpoints less accurate.
@@
In the following, we link the IP to influence functions and provide our insights of IP on why such a naive approximation works well in practice. Let $\nabla v^{\textrm{util}}$$=$$\sum\nolimits_{z_{j}\in V}\nabla_{\hat{\theta}}\ell(z_{j};\hat{\theta})^{\top}$, then influence functions in Eqs. (1) and (2) can be reformulated as $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$ and $\nabla v^{\textrm{util}}\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$. We illustrate the directions of $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ and $\nabla v^{\textrm{util}}$ in Figure 1, with the influence scores roughly representing the angle between the sample gradient and $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ or $\nabla v^{\textrm{util}}$. Remarkably, although influence scores by Eqs. (1) and (2) may differ, they exhibit order-consistency in several scenarios. For instance, if two samples $z_{i}$ and $z_{i^{\prime}}$ have gradients in Region II or IV, then $\mathcal{I}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}^{\textrm{util}}(-z_{i^{\prime}})$ implies $\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i})\gtrless\mathcal{I}_{\textrm{IP}}^{\textrm{util}}(-z_{i^{\prime}})$. However, this order-consistency does not hold in Regions I and III. Fortunately, samples in these regions are jointly recognized by $\mathcal{I}^{\textrm{util}}$ and $\mathcal{I}^{\textrm{util}}_{\textrm{IP}}$ as either beneficial or detrimental, resulting in minimal practical differences. A full analysis is in Appendix A.
@@
In general, IP emerges as a straightforward, efficient, and remarkably effective alternative to influence functions. In non-convex scenarios, approximations of the inverse Hessian matrix may introduce significant errors, rendering $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ ineffective, even incorrect. We substantiate this observation empirically in Section 4. Conversely, while $\nabla v^{\textrm{util}}$ in IP may not precisely align with the optimal direction for enhancing model performance, it remains a reliable indicator for distinguishing between detrimental and beneficial samples, even for non-convex models. Intuitively, if a sample gradient aligns with the gradient from the validation set, it suggests that incorporating this sample contributes to enhancing the model’s utility. Similarly, the negative samples works oppositely.
@@
Figure 2: Illustrating our IP on two synthetic datasets and convex/non-convex models. A-C illustrate a 2D linearly separable synthetic dataset with a subset of detrimental samples bearing incorrect labels, trained using a Logistic Regression model, and D-F demonstrate the similar analysis on a non-linear synthetic half-moon dataset using a Multilayer Perceptron neural network. Specifically, A and D depict training sets with two classes, where detrimental samples are marked with $\times$ and regular samples with $\circ$. B and E show test sets. C and F present influence scores and IP scores by Eqs. (1) and (2), respectively. In the linear case, there is a clear correlation between influence scores and inner product scores, the detrimental samples have both negative influence scores and IP scores. However, in the non-linear case, the influence scores of detrimental samples appear intermixed; but the detrimental samples can be identified via IP.
Enhancement. The simplicity of IP offers opportunities to enhance the generalization of influence functions. In convex cases, the model parameter $\hat{\theta}$ is both optimal and unique. However, in non-convex scenarios, the presence of local minima introduces instability and non-uniqueness into the solution. Typically, ensemble strategies are employed to bolster model generalization [14, 32]. Yet, while employing different models can enhance performance, it also escalates the costs associated with model training and complicates the calculation of influence functions. This arises from the variability in model parameters, necessitating multiple computations of the inverse of each individual Hessian matrix. The introduction of Hessian-free IP circumvents this issue, eliminating the need for costly calculations of Hessian matrices and their inverses. Drawing inspiration from dropout mechanisms, diverse models can be swiftly generated without necessitating model retraining. By computing sample gradients from various models, we propose IP Ensemble that amalgamates IP scores from distinct models. Experiments detailed in Section 5 illustrate the superior performance of the inner product ensemble over other influence function-based methods.
@@ 4 Correctness Verification on Synthetic Data
Here we verify the correctness of our Inner Product (IP) as an influence score surrogate on two synthetic datasets with convex/non-convex models. First, we use Logistic Regression on a linear dataset to demonstrate the effectiveness of IP in a convex optimization scenario. Second, we use a
Figure 3: Directions of gradients of the validation set and detrimental samples in Figure 2. In the linear case, $\alpha$ is $0.66\degree$, and we draw a larger angle for better visualization. For the non-linear case, we use the dimension reduction [7] for visualization, and $\alpha$ is $94.36\degree$.
Multi-Layer Perceptron on a half-moon dataset to validate the performance of IP in a non-convex optimization scenario. [^1] The detailed characteristics of datasets and models used in this paper can be found in Appendix B. In Figure 2, A and B illustrate the training and test sets of a linearly separable dataset, consisting of 150 and 100 samples, respectively. Notably, the training set contains 10 manually generated noisy samples with incorrect labels. Similarly, D-F illustrate a non-linear separable half-moons dataset, consisting of 200 training samples including 20 noisy samples with incorrect labels, and 100 test samples. C and F display the influence score and inner product score for each training sample, calculated according to Eqs. (1) and (2), respectively. Additionally, fairness and robustness analysis is in Appendix C.1.
@@
However, the limitations of influence scores become apparent in the context of non-convex models, as illustrated in Figure 2F. Here, the influence scores of detrimental samples are intermingled with those of normal ones, where $\nabla v^{\textrm{util}}\mathbf{H}^{-1}_{\hat{\theta}}$ is not the optimal direction due to the inaccuracies in approximating the Hessian matrix. Fortunately, the IP score effectively isolates detrimental samples from inliers. Even for this non-linear dataset, $\nabla v^{\textrm{util}}$ remains useful for discerning detrimental samples from beneficial ones, as all detrimental samples exhibit obtuse angles with $\nabla v^{\textrm{util}}$.